n = 68, ¯x=75, and σ = 8. The sample population is normally distributed. Find the 99% interval estimate for µ.

1
Expert's answer
2022-04-13T08:09:08-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{\sigma}{\sqrt{n}},"

where:

  • "\bar{x}=75" : sample mean
  • zthe chosen z-value, for a 99% confidence interval z = 2.576
  • "\sigma=8" sample standard deviation
  • n = 68 sample size.

So,

"CI=75 \pm 2.576\cdot\cfrac{8}{\sqrt{68}}=\\=
75\pm2.50=(72.5, 77.5)."